prove whether this is correct or not :
(n)^2 – (n+1)^2 = n+(n+1)

Grade:8

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given LHS=(n2)-(n+1)2
=n2-(n2+2n+1)
=n2-n2-2n-1
=-2n-1
But RHS=n+n+1=2n+1
As LHS is not equal to RHS the given equation is not correct.
Yugabrat Gogoi
9 years ago
hey Saloni.
 
let in the question n2 – (n+1)2 be lhs.
and n+(n+1) be the rhs.
 
lhs= n2+(n+1)2     {let n be a and(n+1) be b} [formula= a2-b=(a+b)(a-b)]
=[n+(n+1)][n-(n+1)]
=(2n+1)(-1)
=(-2n-1)
 
rhs=[n+(n+1)]
=2n+1
 
therefore, lhs is not equal to rhs.
 
the question is not correct